# DRIFT: Decompose, Retrieve, Illustrate, then Formalize Theorems

### v1

@@ A.2.6 Parametric Retrieval
The results in Table 7 show that external knowledge from retrieval is largely indispensable and cannot be fully replaced by the LLM’s internal knowledge alone. However, we observe a notable difference between the models. DeepSeek-3.1 demonstrates a stronger grasp of the required formal knowledge; for this model, the sub-queries appear to function as a Chain-of-Thought-style prompt, structuring its reasoning process and thereby improving formalization accuracy. This aligns with our earlier finding that DeepSeek-3.1’s zero-shot performance with sufficient sampling (pass@10) nearly saturates to its retrieval-augmented performance, suggesting it can often surface the necessary formal knowledge if properly prompted.
@@ A.2.7 Statistics of Decomposed Sub-queries
To better understand the Decompose module and its behavior, we analyzed the number of sub-queries generated when decomposing informal statements from our three benchmarks: ProofNet, MiniF2F-test and ConNF with different LLMs as decomposer.
@@
| DeepSeek-3.1 | 4.83 | 4.65 | 5.71 | 5.06 |
| Dataset Avg. | 5.69 | 5.21 | 6.42 | 5.77 |
The results, summarized in Table 8, show that different models produce a varying number of sub-queries. GPT-4.1 tends to generate the most detailed decompositions, with an average of 6.27 sub-queries, while DeepSeek-3.1 produces the most concise ones, averaging 5.06. Furthermore, the complexity of the dataset appears to influence the decomposition length. Statements from the ConNF benchmark, which covers frontier mathematical research, consistently required more sub-queries (6.42 on average) across all models, likely reflecting their greater conceptual density compared to the undergraduate-level problems in ProofNet (5.69) and the more self-contained problems in MiniF2F-test (5.21).

### v3

@@ A.2.6 Parametric Retrieval
The results in Table 7 show that external knowledge from retrieval is largely indispensable and cannot be fully replaced by the LLM’s internal knowledge alone. However, we observe a notable difference between the models. DeepSeek-V3.1 demonstrates a stronger grasp of the required formal knowledge; for this model, the sub-queries appear to function as a Chain-of-Thought-style prompt, structuring its reasoning process and thereby improving formalization accuracy. This aligns with our earlier finding that DeepSeek-V3.1’s zero-shot performance with sufficient sampling (pass@10) nearly saturates to its retrieval-augmented performance, suggesting it can often surface the necessary formal knowledge if properly prompted.
@@ A.2.7 Statistics of Decomposed Sub-queries
To better understand the Decompose module and its behavior, we analyzed the number of sub-queries generated when decomposing informal statements from our three benchmarks: ProofNet, MiniF2F-test, and ConNF with different LLMs as decomposers.
@@
| DeepSeek-V3.1 | 4.83 | 4.65 | 5.71 | 5.06 |
| Benchmark Avg. | 5.69 | 5.21 | 6.42 | 5.77 |
The results, summarized in Table 8, show that different models produce a varying number of sub-queries. GPT-4.1 tends to generate the most detailed decompositions, with an average of 6.27 sub-queries, while DeepSeek-V3.1 produces the most concise ones, averaging 5.06. Furthermore, the complexity of the benchmark appears to influence the decomposition length. Statements from the ConNF benchmark, which covers frontier mathematical research, consistently required more sub-queries (6.42 on average) across all models, likely reflecting their greater conceptual density compared to the undergraduate-level problems in ProofNet (5.69) and the more self-contained problems in MiniF2F-test (5.21).